OpenStudy (anonymous):

A rock falls off of a 2.00 m high wall. When it hits the ground, the ground compresses a distance of 1.00 cm as it comes to a stop. What was the acceleration of the rock as it came to a stop after contacting the ground? Assume the motion during the stop is with constant acceleration.

OpenStudy (anonymous):

If the acceleration was constant it s 0 .

OpenStudy (anonymous):

@E.ali No. Zero is not part of the answers.

OpenStudy (anonymous):

Well .... Then you cant ask the question good ... Can you ask better ?

OpenStudy (fifciol):

Since a is constant \[a=\frac{ v }{ t }\Rightarrow a^2=\frac{v^2}{t^2}\] \[d=\frac{ at^2 }{2 } \Rightarrow t^2=\frac{2d}{a}\] Conservation of mechanical energy \[mgh=\frac{ mv^2 }{ 2 } \Rightarrow v^2=2gh\]so \[a^2=\frac{ 2gha }{ 2d} \Rightarrow a=\frac{ gh }{ d }\]

OpenStudy (fifciol):

about 1962 m/s^2

OpenStudy (anonymous):

@Fifciol do you mean19.62 m upward?

OpenStudy (fifciol):

convert cm to m you'll get then 1962 m/s^2 pointing upward

OpenStudy (anonymous):

@Fifciol Oh, I see. Thanks.

OpenStudy (fifciol):

yw

OpenStudy (anonymous):

Ok ! I didnt know your question!